You ask a friend to think of a number from 4 to 10. What is the probability that his number will be a multiple of three?


1
Expert's answer
2021-05-04T12:42:31-0400

S= number from 4 to 10 = {4,5,6,7,8,9,10}

A = number will be a multiple of three = {6,9}

Probability = n(A)/n(S) = 2/7